OpenStudy (agentnao):

What is the sum of an 8-term geometric sequence if the first term is 10 and the last term is 781,250?

OpenStudy (anonymous):

\[2\times5^1 + 2\times5^2 + 2\times5^3+2\times5^4+2\times5^5+2\times5^6+2\times5^7+2\times5^8 = ?\]

OpenStudy (anonymous):

976560

OpenStudy (agentnao):

Could anyone please tell me how to find r? Because I have the formula, but I don't konw how to find r (the ratio)

OpenStudy (anonymous):

I thought all you needed was the sum

OpenStudy (agentnao):

Well, yes, that's all the problem asks for, but using the given formula, I can't solve it without knowing what r is. I have to show my work for the answer but I can't figure out how to find r.

OpenStudy (anonymous):

What formula do you have?

OpenStudy (agentnao):

Sn = a1 - anr/(1-r) The an equals the last term and the r is, of course, the ratio.

OpenStudy (anonymous):

r = last term/previous term = 2*5^2/2*5^1 = 5; r = 5

OpenStudy (anonymous):

let's see if we get the same answer if we apply that

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

For some reason, I got 976550, instead of 976560, I'm off by like 10

OpenStudy (anonymous):

\[\frac{(10-781250)5}{1-5}\]

OpenStudy (agentnao):

Wait. r = last term?

OpenStudy (anonymous):

r = 5, I posted that above :/

OpenStudy (agentnao):

r is supposed to be the ratio between the terms in the sequence, not the last term. And yes, I saw that.

OpenStudy (anonymous):

r = last term/previous term = 2*5^2/2*5^1 = 5; r = 5

OpenStudy (anonymous):

I never said r was the last term

OpenStudy (agentnao):

Ohhhh. Oh goodness, I'm sorry. I misread what you wrote. Seeing division problems written horizontally confuses me if I'm not paying strict attention.

OpenStudy (anonymous):

What I posted was actually what happens when you put r into your formula: \[\frac{(10-781250)5}{1-5} \]

OpenStudy (anonymous):

Ohhhhhhh

OpenStudy (agentnao):

But how did you get 5?

OpenStudy (anonymous):

I know what happened

OpenStudy (agentnao):

Would you mind joining me in a whiteboard room so you can show me what you're doing?

OpenStudy (agentnao):

http://www.twiddla.com/594690 <-- that's the ling

OpenStudy (anonymous):

Here's the formula: \[\frac{10 - (781250*5)}{1-5}\]

OpenStudy (agentnao):

*link

OpenStudy (agentnao):

I have that same formula, but i still have the variable r in place of your 5s. I don't understand how you came to that.

OpenStudy (agentnao):

If you don't want to do the room thing, that's fine too. I hope you didn't think that was weird or creepy because that's definitely not how I intended it to seem.

OpenStudy (anonymous):

I came to the room. You don't see my post?

OpenStudy (anonymous):

r = 5, c'mon....Why are you not understanding this? In order to get the sum, you have to substitute everything, including r

OpenStudy (anonymous):

Do you see my twiddla posts or not?

OpenStudy (agentnao):

Yes!

OpenStudy (anonymous):

Does it make any sense now?

OpenStudy (agentnao):

How did you know what the binomial was? That it was 2 x 5^1